OpenStudy (anonymous):

A 20 foot ladder is placed against a building to reach a window that is 16 feet above the ground. How many feet away from the building is the bottom of the ladder? (Points : 2) 3 4 12 15 @jim_thompson5910 i put 12

OpenStudy (anonymous):

you are correct it is 12

OpenStudy (anonymous):

If you draw it out, you'll notice that you have a right triangle with the ladder being the hypotenuse. That means that the distance from the building squared plus the height above the ground squared equals the length of the ladder squared: \[16^{2} + f^{2} = 20^{2}\]\[256 + f^{2} = 400\]\[f^{2} = 400 - 256\]\[f^{2} = 144\]\[f = \pm 12\] -12 doesn't make any sense here so the answer is 12.